OpenStudy (anonymous):

The measure of each interior angle of a regular polygon is 170. Find the number of sides in the polygon. PLEASE HELP!

OpenStudy (kagıtucak):

inner angle is 170 then outer angle is 180-170=10 sum of the outer angle 360 so 360/n=10 n=36

OpenStudy (anonymous):

@kagıtucak what is the formula you used?

OpenStudy (kagıtucak):

i didnt use a formula just apply basic things

OpenStudy (anonymous):

becausei know (n-2) x 180 is for finding the sum of the interior angles of a polygon

OpenStudy (kagıtucak):

for exapmle think equilateral triangle its inner angle 60 so outer 120 360/n=120 n=3 and obviously triangle has 3 sides

OpenStudy (anonymous):

how did you know to do 360/10 is what im confused about

OpenStudy (kagıtucak):

sum of outer angle is always 360 for every close shape

OpenStudy (anonymous):

yeah i understand how you got the numbers but how did you know to divide?

OpenStudy (kagıtucak):

and our inner angles are equal for each side

OpenStudy (anonymous):

but why did you divide?

OpenStudy (kagıtucak):

we know one outer angle and all outer angles are equal and their sum is 360 so how many outer(10) have to come together to make 360

OpenStudy (anonymous):

oh kk thanks!!!!